price of coffee is raised to $1.80, then 120 cups of tea are sold. 12. Assume the quantity demanded of peanut butter fell byCan someone explain 13 and 14?
Answer #1

13. Cross price elasticity of demand = % change in the quantity demanded of peanut butter/% change in the price of strawberry jam = -2/10 = -0.2

As the cross-price elasticity of demand is negative, the two goods are complements. (For substitutes, the cross-price elasticity of demand is positive)

14. Initial revenue from cheeseburgers = (Pc x Qc) = $1000

Initial revenue from soft drinks = (Ps x Qs) = $600

Total revenue = $1000 + $600 = $1600

After changes in the prices,

Price of Cheeseburgers = (1+0.10) X Pc = 1.1Pc

Quantity demanded of Cheeseburgers = (1-0.05)Qc = 0.95Qc

Revenue from Cheeseburgers= 1.1Pc x 0.95Qc = 1.045 (Pc x Qc) = 1.045 x $1000 = $1045

Price of soft drinks = Ps (remains the same)

Quantity demanded of Soft drinks = (1-0.10) x Qs = 0.90Qs

Revenue from Soft drinks = Ps x 0.90Qs = 0.90 (Ps x Qs) = 0.90 x $600 = $540

Total Revenue = $1045 + $540 = $1585

Therefore, the total revenue decreased due to a 10% increase in the price of cheeseburgers.